determine whether each of the following ordered pairs are solutions to the linear equations given
(-3,-8):y=5x+7?
(3,4):5x-2y=6

Answer by rfer(16322)

y=5*-3+7
5=-15+7
y=-8
number one Yes
--------------------
5x-2y=6
5x=2y+6
5x=2*4+6
5x=8+6
5x=14
x=14/5
x=2.8
number two No